For the calculation of caloric temperature please refer [3] (page 827). 


1
Expert's answer
2021-09-24T02:13:54-0400

We use the equation below for the calculation of caloric temperature:




We have °API of hot fluid=76°; Therefore Kc = 1; Fc = 0.455

Caloric temperature for hot fluid becomes:

120+0.455×(160-120) =138.2°F

Caloric temperature of the cold fluid becomes:

=75+0.455×(120-75) =95.475°F